OpenStudy (anonymous):

what am i doing wrong? A copper pipe, used for plumbing from the water main to a house, is 4 meters long and has a circular cross-section 2.8 cm in diameter. There is a 2 cm diameter hole in the center of the cross-section for the entire length of the rod. What is the volume of copper in the rod in cubic inches? The density of copper is 559lb_m/ft^3 . What is the mass of the copper pipe in pounds? in caculating the mass of copper what am i doing wrong? pi(1.4)^2*500-pi(1)^2*500= 1507.9 1507.9 * 01974=29.76 lbs

OpenStudy (bahrom7893):

hold on, i'm trying to understand the question

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

it's 400 centimeters u're multiplying by 500

OpenStudy (bahrom7893):

Does that answer your question?

OpenStudy (anonymous):

can you help me with finding the cubic inches

OpenStudy (anonymous):

volume of cooper in cubic inches

OpenStudy (bahrom7893):

Well is this answer right in centimeters: 1507.9 cm^3 ? I don't think so, because u multiplied by 500, when u were supposed to multiply by 400

OpenStudy (anonymous):

ocond part of the question, i now understand. can you help me with first part pls

OpenStudy (anonymous):

second

OpenStudy (bahrom7893):

which part do u need help with? First?

OpenStudy (anonymous):

yes

OpenStudy (bahrom7893):

No that's the thing the 2nd part depends on the first and the first part is wrong. Anyway..

OpenStudy (anonymous):

i was able to find the mass pounds without doing the cooper

OpenStudy (anonymous):

I believe the diameter was given in the question and the cross section is area of the circle, which means you should be using Area = Pi * radius ^2

OpenStudy (bahrom7893):

If we use the correct numbers: pi(1.4 cm)^2*400 cm-pi(1 cm)^2*400 cm = pi(1.4 * 0.39 in)^2*400 cm-pi(1 * 0.39 in)^2*400 * 0.39 in

OpenStudy (bahrom7893):

One cm is 0.39 inches, so all you gotta do is just convert your centimeters into inches in the formula and that's it

OpenStudy (bahrom7893):

woops forgot to multiply by 0.39 at one point pi(1.4 * 0.39 in)^2*400 * 0.39 in-pi(1 * 0.39 in)^2*400 * 0.39 in =

OpenStudy (bahrom7893):

71.568in^3 is the volume

OpenStudy (anonymous):

thanks